Can you help me with this equation please? I have to write this equation is standard form: y-4=5/3(x+3)
Answer by stanbon(75887)

write this equation is standard form:
y-4= (5/3)(x+3)
Multiply thru by 3 to get:
3y - 12 = 5x +15
----
Standard Form: 5x - 3y = -27
==============================
Cheers,
Stan H.
==============